OpenStudy (anonymous):

Find the simple interest and the total amount due on the loan. Principal Rate Time $18,000 2 1/2% 3 months

OpenStudy (nubeer):

well it will be done by same method as in the previous question u postd

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i tried its confusing

OpenStudy (anonymous):

i plugged in same as the other one i get it but than i dont

OpenStudy (anonymous):

help

OpenStudy (nubeer):

ok its like (18000)(5/2)(3/12)(100) thats it

OpenStudy (anonymous):

ok

OpenStudy (nubeer):

did u tried like this..? still wanna c how u tried

OpenStudy (anonymous):

so 1125000

OpenStudy (nubeer):

yes :)

OpenStudy (nubeer):

just let me know if i get the wrong answer.

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

im getting frustrated:( i cant do this

OpenStudy (nubeer):

look i will tell u in detail intrest = (18000)(2.5/100)(3/12)(100) intrest = 11250

OpenStudy (nubeer):

total amount = 18000+11250 =29250